The perimeter of a rectangle is 28 cm. If the length were subtracted by 2 and the width added by 2, the rectangle would have been a square. What are the dimensions of the rectangle?
Answer by reviewermath(1029)

Let x = width and y = length
The perimeter is 28 cm so 2x + 2y = 28 or x + y = 14.
Also, y - 2 = x + 2 or x - y = -4
Add the equations x + y = 14 and x - y = -4, the result is
2x = 10
x = 5
so y = 9
Answer: length = 9 cm and width = 5 cm.